Find the directional derivative of \\(Ï•=xy^2+yz^2+xyz\\) at the point (2, -1, 1) in the direction of the vector \\(A=2i+4j-k\\).
1
Expert's answer
2020-11-17T16:18:45-0500

"f(x, y, z)=xy^2 +yz^2+xyz, M(2, -1, 1)"


"\dfrac{\partial f}{\partial x}=y^2+yz,"


"\dfrac{\partial f}{\partial y}=2xy+z^2+xz,"

"\dfrac{\partial f}{\partial z}=2yz+xy"

"\dfrac{\partial f(M)}{\partial x}=(-1)^2+(-1)(1)=0,"

"\dfrac{\partial f(M)}{\partial y}=2(2)(-1)+(1)^2+2(1)=-1,"

"\dfrac{\partial f(M)}{\partial z}=2(-1)(1)+2(-1)=-4"

Given "\vec{u}=2\vec{i}+4\vec{j}-\vec{k}"

"|\vec{u}|=\sqrt{(2)^2+(4)^2+(-1)^2}=\sqrt{21}"

"\vec{l}=\dfrac{2}{\sqrt{21}}\vec{i}+\dfrac{4}{\sqrt{21}}\vec{j}-\dfrac{1}{\sqrt{21}}\vec{k}"

"\dfrac{\partial f}{\partial l}=0\cdot\dfrac{2}{\sqrt{21}}+(-1)\cdot\dfrac{4}{\sqrt{21}}+(-4)\cdot(-\dfrac{1}{\sqrt{21}})=0"

The directional derivative 


"\dfrac{\partial f}{\partial l}=0"